OpenStudy (anonymous):

Find the solutions of the equation. 1/2x^2-x+5=0

OpenStudy (anonymous):

easier to get rid of the fraction and start with \[x^2-2x+10=0\]

OpenStudy (anonymous):

the solutions will not be real, they will be complex numbers. is that ok?

OpenStudy (waleed_imtiaz):

Multiply whole equation by 2..... u will get.... x^2-2x+10=0 and now u can use quadratic equation or make factors.......

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

 This is what my answers look like

OpenStudy (anonymous):

this is not going to factor with real numbers you can use the quadratic formula if you like or ...

OpenStudy (anonymous):

yes, but since \(\sqrt{9}=3\) you get \(1\pm3i\)

OpenStudy (anonymous):

Okay! thank you.